### v1

# Alternating Perception-Reasoning for Hallucination-Resistant Video Understanding

Abstract
Sufficient visual perception is the foundation of video reasoning. Nevertheless, existing Video Reasoning LLMs suffer from perception shortcuts, relying on a flawed single-step perception paradigm. This paradigm describes the video and then conducts reasoning, which runs the risk of insufficient evidence and emergent hallucinations. To address these issues, we introduce a new framework that integrates a loop-based paradigm with an anti-hallucination reward. First, to address the insufficient evidence, we introduce the Perception Loop Reasoning (PLR) paradigm. Instead of describing the video at once, each loop requires the model to describe a video segment with precise timestamps, analyze this segment, and decide the next action. Second, for the risk of hallucinations, the Factual-Aware Evaluator (FAE) evaluates each perception result as a reliable anti-hallucination reward. This reward encourages the model to provide sufficient and precise video evidence. Our FAE, which performs comparably to GPT-4o, is tuned on our AnetHallu-117K, a large-scale hallucination judgment preference dataset. Extensive experiments show that our Video-PLR achieves the state-of-the-art in both 3B and 7B parameter scales and has the best data efficiency. Our code, models, and datasets are released on: https://github.com/BoweiPu/VideoPLR.
Figure 1: Overview of VideoPLR. (a) Compared with reasoning paradigms. VideoPLR performs iterative perception and reasoning, and each perception execution targets the perception results of the previous reasoning. During training, it is allowed to individually judge whether hallucinations occur in each perception result and provide rewards. (b) Examples of VideoPLR. The PLR paradigm perceives partial video clips multiple times, achieving repeated perception and providing correct answers.(c) Benchmark performance: Comparative results on 7 benchmarks highlight the video reasoning capabilities of our model, at both 3B and 7B parameter scales.
1 Introduction
Video Large Language Models (VideoLLMs) [llava-video, liLLaVAOneVisionEasyVisual2024, wangInternVideo2ScalingFoundation2024a] demonstrate remarkable potential in traditional video understanding. Building on this success, researchers are increasingly turning their attention to complex video tasks. Inspired by the Think-then-Answer paradigm [guo2025deepseek], many VideoLLMs [wang2025videorft, zhang2025tinyllava, feng2025video] tackle complex problems by generating intermediate reasoning steps. This reasoning is based on video perception: the accurate description of visual information in the video. Thus, insufficient perception or hallucinations often lead to flawed or erroneous reasoning.
The previous paradigm is to describe the video first and then reason. This paradigm has fundamental flaws. As shown in Fig. 1(b), the video description in the Previous Reasoning overlooks the video segment of adding bean sprouts, losing the key evidence. Also, this reasoning is inaccurate and exhibits hallucinations. This example shows that the single-step perception faces two challenges: (1) Partial and Incomplete Perception: Single perception often analyzes a part of the video, lacking a mechanism to assess the sufficiency of the evidence. This leads to the loss of key evidence. (2) Risk of Hallucinatory Perception: The entire video contains rich and dynamic visual information, making it difficult to assess the accuracy of the overall description. This leads to a loss of reliable hallucination supervision.
Based on the analysis of single-step perception, we argue that a better paradigm should enable the model to repeatedly perceive different video segments and provides reliable hallucination supervision during training. The number of perceptions should be decided automatically by the model itself. Thus, we introduce Perception Loop Reasoning (PLR) paradigm, which reformulates this single-step perception into a multi-step paradigm, as shown in Fig. 1(a). The core idea of PLR is to explicitly decouple Perception and Reasoning into repeating steps. Different from traditional video describing, PLR requires the model to provide timestamps and corresponding descriptions. After each perception, the model will analyze this segment and decide whether to output an answer or continue to perceive. In this way, our PLR paradigm ensures that the completeness of the reasoning is grounded in accurate and sufficient perceptual evidence gathered iteratively. Fig. 1(b) demonstrates multi-step and precise perception.
Realizing this multi-step perception is challenging, as current models lack the ability to re-perceive, and no dataset exists to support it. Following the PLR paradigm, we construct VideoPLR-14K, a novel dataset designed to bootstrap the model to generate responses in the PLR format, encompassing both sequential causal reasoning and non-sequential segment ordering. Furthermore, anti-hallucination is the key to precise perception, which is the foundation of reasoning. However, previous models, including Open-source VideoLLMs [wang2024videohallucer, li2025vidhalluc] and vision-language models [zhai2023sigmoid], cannot provide reliable factual evaluation. To fill this gap, we introduce a Factual-Aware Evaluator (FAE), which is meticulously trained on our newly constructed AnetHallu-117K dataset. The ANetHallu-117K is constructed from the human-annotated ANet [caba2015activitynet] and contains 5 types of hallucination and hybrid CoT binary classification preference data, which can reduce discriminative bias. FAE is comparable to GPT-4o in evaluating text-video consistency to serve as a reliable anti-hallucination reward for RL training. The segmented approach of PLR is also more reliable for hallucination evaluation. Specifically, their constrained visual context simplifies grounding the descriptions, making anti-hallucination more readily identifiable.
These two efforts jointly constitute our final model, Video-PLR via reinforcement learning [guo2025deepseek], achieving high data efficiency. For the RL stage, we utilize only 6K of mixed data, including five types of tasks. Our models (3B and 7B) achieve SOTA results on multiple benchmarks, as illustrated in Fig. 1(c). Compared with previous models, Video-PLR integrates accurate timestamps directly into the reasoning text, enabling a unified consideration of temporal grounding. It is worth mentioning that by simply adding 0.5K of Video Temporal Grounding data, we outperform models trained with more than 5 times the amount of data. PLR not only emphasizes the necessity of sufficient evidence but also takes into account time perception, visual content perception, and logical reasoning, aiming to build a more advanced video reasoning paradigm.
The contributions of this paper are as follows:
  • We propose Perception Loop Reasoning (PLR), a novel video reasoning paradigm that addresses the perception shortcut. PLR reformulates video reasoning from a single-step process into a multi-step perception and reflection loop, where the model iteratively perceives segments and analyzes them before concluding.
  • To enable fine-grained hallucination evaluation within the loop, we introduce a Factual-Aware Evaluator (FAE) and its corresponding large-scale preference dataset, AnetHallu-117K. The FAE reaches the level of GPT-4o in hallucination benchmarks and generates anti-hallucination rewards in RL training.
  • We construct VideoPLR-14K, a dataset designed to bootstrap the PLR paradigm. Our resulting model, Video-PLR, trained via RL using FAE as an anti-hallucination reward, achieves SOTA performance both on 3B and 7B parameter scales with data efficiency.
2 Related Work
2.1 Multimodal Chain-Of-Thought
Multimodal Chain-of-Thought (MCoT) simulates human reasoning by decomposing complex problems into intermediate steps, thereby enhancing transparency and performance. LLaVA-CoT [xu2024llava] is the first to implement a multimodal O1-like [openai2024o1] reasoning approach. Similarly, Video-of-Thought [feivideo] proposes a comprehensive five-stage framework. VideoEspresso [han2025videoespresso] and VideoCoT [wang2024videocot] achieve MCoT by invoking external tools as a multi-stage approach. Chain-of-Frames [ghazanfari2025chain] further integrates this capability into the model’s internal thinking process, enabling it to describe keyframe content. However, single-frame descriptions are static, making it difficult to perform dynamic analysis across frames. We argue that the video modality necessitates CoT because videos contain vast amounts of information, requiring the model to identify crucial details while filtering out noise. The model’s process of progressively identifying key information is inherently dynamic, and our designed perception and analysis loop supports this process.
2.2 Reasoning VideoLLMs
VideoR1 [feng2025video] first achieves R1-like [guo2025deepseek] video reasoning and utilizes GRPO [shao2024deepseekmath] for training. However, its cold-start data is sampled from Qwen2.5VL 72B [bai2025qwen25vltechnicalreport]. This strategy overlooks the fact that even models with larger parameter counts lack video reasoning training. To enhance video perception, VideoR1 and DeepVideo-R1 [park2025deepvideo] enhance the model’s perception by modifying the standard GRPO. VideoRFT [wang2025videorft] advances this further by restructuring the CoT into a fixed pipeline: task, perception, reasoning, and conclusion. It utilizes SigLip [tschannen2025siglip] during the perception stage to mitigate hallucinations. TimeR1 [wang2025time] and VideoChatR1 [li2025videochat] place a greater emphasis on temporal perception and often lack consideration for general-purpose tasks. VideoChatR1.5 [yan2025videochat] achieves tool-use for video trimming based on temporal reasoning. ARC-HunyuanVideo [ge2025archunyuanvideo7bstructuredvideocomprehension] can generate dense video descriptions with timestamps, but lacks the ability to reason within these segments, and also does not address the hallucination problem. In contrast, our reasoning paradigm integrates temporal perception and execution, anti-hallucination training, and cyclical evidence gathering, representing a more advanced approach to reasoning training.
2.3 Video Hallucination
Longer reasoning chains reduce attention to visual input, thereby leading to hallucinations [liu2025more]. Consequently, hallucination control is crucial for video models. VideoRFT [wang2025videorft] utilizes SigLip [tschannen2025siglip]. Nevertheless, SigLip is not optimally suited for video. Utilizing a VideoLLM as a judge requires the model itself to be low-hallucinating and unbiased. However, the current models cannot achieve this. VidHalluc [li2025vidhalluc] evaluates hallucinations from three key dimensions, while VideoHallucer [wang2024videohallucer] focuses on the bias of judging the correctness of descriptions. To address this, we provide a controllable data generation pipeline for hallucination pairs and introduce a video hallucination judgement model. Our primary objective is to supply a reliable anti-hallucination reward.
3 Factual-Aware Evaluator
Figure 2: The pipeline for hallucination preference data and cold start data. (a) The figure above shows the construction method of hallucination data. By re-annotation, five types of hallucinated captions are generated. Then, Qwen 2.5VL-7B is used to automatically generate CoT data, resulting in the Factual-Aware Evaluator. (b) The figure below shows the pipeline for building cold start data following the PLR paradigm. It mainly demonstrates the method of reannotating NextQA, using carefully designed prompts to enable the model to generate inference data containing perceptual loops.
(a) Perception Number
(b) RL data type
Figure 3: Statistics of the post-training dataset
(a) Negative Captions
(b) Positive Captions
Figure 4: Comparison of caption lengths in ANetHallu-117K
To provide reliable consistency between textual descriptions and visual content, Factual-Aware Evaluator is trained on AnetHallu-117K, which is based on human-annotated ActivityNet (ANet) [caba2015activitynet]. In this section, we describe in detail the caption construction pipeline and the preference training pipeline, as shown at the top of Fig. 2. The key to this pipeline is to eliminate the bias of language and strengthen the visual judgment of the model.
3.1 Re-annotate Dense Caption
ANet contains video content from various scenarios; however, its original annotations are coarse-grained. It is impossible to obtain hallucinated caption pairs from such annotations. Thus, we first generate high-fidelity dense descriptions as ground-truth, $C_{gt}$, for each video segment $t_{1},t_{2}]$ utilizing Qwen2.5-VL 32B based on original annotations. The original human annotations guided the model’s video description process, achieving near-realistic dense descriptions.
3.2 Low-Bias Judgement
This part includes the pipeline for generating low language bias caption pairs and enhancing perception using CoT.
First, the requirement of low language bias means that it is difficult to distinguish whether it is a hallucination caption through language. We devised a two-stage Hallucinate-then-Correct pipeline to maintain the language style and sentence length consistency within a caption pair. A broad range of hallucination types brings better robustness. There are five hallucination types: (1) Attribute Modification, (2) Quantity Modification, (3) Action Substitution, (4) Detail Conflation, and (5) Temporal Reordering. By providing the model with $C_{gt}$ and the hallucination definition, negative captions of the specified hallucination types can be obtained. For the correction stage, provide the model with both $C_{neg}$ and $C_{gt}$ simultaneously. The positive captions, $C_{pos}$, generated by the model serve as a direct fix for the specific error present in $C_{neg}$. Thus, raw negative and positive caption pairs are collated. Fig. 4 presents the caption pairs that have similar language styles and text lengths.
Because of the nature of LLMs, captions exhibit significant differences in vocabulary distribution. It may cause the vulnerability to be discovered during model training and result in the model exploiting this vulnerability. Text Bias Elimination step removes these captions containing the significant bias words, which calculates a bias score for each caption using the biased vocabulary sets $\mathcal{W}_{pos}$ and $\mathcal{W}_{neg}$. This biased vocabulary set comes from the word frequency ratio. Relative frequency ratio $R_{pos/neg}$ is computed for each word $w$ in the positive and negative sample sets. Specifically, for words in negative caption sets, $R_{neg}$ is defined as:
$R_{neg}(w)$ $=\frac{f_{neg}(w)}{\max(f_{pos}(w),1)},$ (1)
where $f(w)$ is the total frequency of word $w$ in the corresponding set. The biased vocabulary set is defined as the top $N$ words with the highest $R_{pos/neg}(w)$ in each caption set. We assign penalty scores $P(w)$ with these biased words.
$P_{neg}(w)=\frac{R_{neg}(w)}{\max_{w^{\prime}\in\mathcal{W}_{neg}}R_{neg}(w^{\prime})}.$ (2)
Subsequently, we obtain the total preference score $S(c_{neg})$ for a given caption $c_{neg}$ by summing the penalties of all its constituent bias words:
$S(c_{neg})=\sum_{w\in W(c_{neg})\cap\mathcal{W}_{neg}}P_{neg}(w),$ (3)
where $W(c_{neg})$ represents the words in the caption $c_{neg}$. We apply the same process to handle positive captions. We finally execute it 15 times, deleting 2% of the data each time, and setting the bias vocabulary set size $N$ as 30.
Despite word-level frequency bias being eliminated well, the model might still exploit subtle linguistic differences (e.g., grammatical fluency) between $C_{pos}$ and $C_{neg}$ to exploit spurious correlations. Thus, we argue that CoT Enhance Perception, which first mandates the model to generate video descriptions. This CoT allows the model to rely on visual details, identify errors, and subsequently arrive at a conclusion.
For a given video clip $V_{[t_{1},t_{2}]}$ and a caption $C_{test}$, we prompt the model to perform a perceptual re-reasoning process: first, generate a brief factual description $R_{desc}$ of $V_{[t_{1},t_{2}]}$, and then, judge whether $C_{test}$ is correct or incorrect. The final data tuple is formatted as $(V_{[t_{1},t_{2}]},C_{test},R_{desc},A)$, where $A$ is the final binary judgment. We select responses where the model accuracy is neither 0% nor 100% as preference data for training.
3.3 Hybrid Preference Training
In the training stage, we construct a mixed training dataset $\mathcal{D}$, which contains multiple types of preference pairs $(x,y_{w},y_{l})$, where $x$ is the input (video-caption pair). The first type is Answer Accuracy (Direct Answer vs. Direct Answer), which prioritizes rapid response. The second type is Reasoning Accuracy (CoT vs. CoT), compelling the model to engage in perceptual reasoning. This mixed data strategy ensures the model not only judges the correctness of the caption but also learns why, thereby achieving robust and efficient video factual evaluation.
Odds Ratio Preference Optimization (ORPO) [hong2024orpo] is employed for training. The total loss function $\mathcal{L}_{ORPO}$ for ORPO is defined as a weighted sum of the SFT loss and the Odds Ratio (OR) preference loss:
$\mathcal{L}_{ORPO}=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\mathcal{L}_{SFT}+\lambda\cdot\mathcal{L}_{OR}\right],$ (4)
where the $\mathcal{L}_{SFT}$ term, $\mathcal{L}_{SFT}=-\log P_{\theta}(y_{w}|x)$, performs standard supervised fine-tuning exclusively on the winning response $y_{w}$.
The $\mathcal{L}_{OR}$ term is the preference loss, which penalizes the model’s odds of ranking $y_{l}$ higher than $y_{w}$:
$\mathcal{L}_{OR}=-\log\sigma\left(\log\frac{\text{odds}_{\theta}(y_{w}|x)}{\text{odds}_{\theta}(y_{l}|x)}\right).$ (5)
All of these prompts and more details about the dataset can be found in the supplementary materials.
4 Perception Loop Reasoning Post-Training
In this section, the Perception Loop Reasoning paradigm is introduced in detail. Then, following this paradigm, we construct SFT data as cold start. In the RL stage, we specifically incorporated an anti-hallucination reward mechanism to encourage the model to generate reasoning chains with accurate facts.
4.1 PLR Design
The core concept of PLR involves iteratively describing designated video segments while dynamically evaluating the perceptual process. The PLR-style reasoning is structured into three phases: initial analysis, iterative perception&reasoning, and summarization&conclusion, depicted in the Fig. 1(a). In the first stage, the model should analyze the question independently and determine which visual content may be crucial. The key to PLR lies in the second stage. After each perception, it is necessary to further analyze the relationship between the visual content of the perception and the problem and determine whether it is sufficient to solve the problem. If the problem can be solved, it will directly enter the third stage; otherwise, it will re-execute the perception action, formulate a new perception goal, and output the next perception result. Therefore, this loop is determined autonomously by the model. Each perception follows a specific format, including timestamps and descriptions of the video clip. This allows for independent evaluation of every perception outcome, yielding more granular assessment results compared to full video descriptions. We employ FAE for hallucination evaluation and encourage the model to generate more perception results. Finally, in the third stage, the model synthesizes the perceptual and reasoning content to derive the answer to the question. This special reasoning path requires cold-start data, and next, we will explain how to construct such SFT data from traditional datasets.
4.2 Cold-Start Data Construction
Video perception is divided into sequential and disordered ones on the time axis. To accommodate the reasoning for these two sequences, we have covered two types of tasks. For the sequential part, we select NextQA [xiao2021next], which contains causal and temporal action reasoning. For the disordered part, we set event ordering based on ActivityNet [caba2015activitynet] named ReOrdering. The challenge is the lack of timestamp and reasoning annotations within NextQA. Thus, we mainly introduce how to construct the CoT with timestamps of NextQA, as shown at the bottom of Fig. 2(b).
Entity Name Extraction: VideoQA suffers from language inconsistency, where synonymous expressions like toy dog and plush toy cause referential ambiguity. To mitigate this, we employ an LLM to extract a normalized entity list $E$ from the question $Q$. This standardizes the terminology and unifies the focus for subsequent steps.
Frame Level Captioning: The objective of this step is to acquire detailed visual descriptions for each frame. To maintain descriptive consistency, when captioning frame $F_{t}$, we feed the VLM with the current frame $F_{t}$, the preceding frame $F_{t-1}$, the normalized entity list $E$, and the previous caption $C_{t-1}$. This prompts the VLM to generate dense descriptions focusing on the relevant entities, their behaviors, and the inter-frame changes.
Raw CoT and Refine Logic: To obtain the PLR-style CoT from the original description, we divide it into a two-stage optimization. The first stage focuses on generating coherent multi-paragraph descriptions. The second stage focuses on adding logic and self-plan.
For the ReOrdering task, the dense captions in sec 3.2 serve as available evidence. For a video containing multiple events, rearrange the short video descriptions into a ReOrdering task, where the answers are the events arranged in chronological order. For NextQA, we have the basic annotation for each frame, but lack a coherent description of video segments. We feed all descriptions, their timestamps, the question, and the answer into an LLM, tasking it with aggregating these inputs into a structured event list $V=\{(\text{start},\text{end},\text{desc})\}$ within the thinking. Crucially, this step incorporates the FAE to rigorously filter the description of each segment. This process eliminates all inaccurate or fabricated events, ensuring factual accuracy. Although we obtain the Raw PLR-style CoT, there is a lack of logic.
In the second stage, we use carefully crafted prompts to enable the LLM to perform this automated data rewriting. Additionally, at this stage, training samples that do not conform to logic should be removed. Finally, we select 14K high-quality data from the original annotations for cold start, and the perceived numbers are diverse, as shown in Fig. 3. All prompts and examples of data will be provided in the supplementary materials.
4.3 Training with GRPO and Reward Design
The Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] process begins by sampling $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$ for a given query $q$ and evaluating their rewards $\{r_{i}\}_{i=1}^{K}$. The rewards are then normalized to compute the advantage $A_{i}$ for each response:
$\small A_{i}=\frac{r_{i}-{\texttt{mean}(\{r_{1},\ldots,r_{K}\})}}{{\texttt{std}(\{r_{1},\ldots,r_{K}\})}},$ (6)
where mean and std are calculated over the $K$ rewards.
The policy $\pi_{\theta}$ is then optimized by maximizing the following objective function:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{[q,\{o_{i}\}]}\frac{1}{K}\sum_{i=1}^{K}\Big[\min\Big(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}}A_{i},$
$\quad\text{clip}\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}},1-\!\epsilon,1+\epsilon\right)A_{i}\Big)-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Big].$ (7)
Here, $\pi_{\theta_{\text{old}}}$ is the policy from the previous iteration, $\pi_{\text{ref}}$ is the reference policy, $\epsilon$ is the clipping coefficient for the policy ratio, and $\beta$ controls the KL divergence penalty.
The key to GRPO is the reward design. Our goal is to ensure the model responds accurately while minimizing hallucinations. The output of the model will be defined as $y$.
Think Format Reward: $y$ must meet the formatting :
<think>...</think><answer>...</answer>. Therefore, format rewards, $R_{f_{t}}$ are defined as:
$R_{f_{t}}(y)=\begin{cases}1,&\text{if }y\text{ matches the think format,}\\ 0,&\text{otherwise.}\end{cases}$ (8)
Evidence Format Reward The model’s perception results are wrapped in a special format, which is required to be <start="t1",end="t2",desc="description">. Therefore, the evidence format reward function is:
$R_{f_{e}}(y)=\begin{cases}1,&\text{if }y\text{ matches the evidence format},\\ 0,&\text{otherwise.}\end{cases}$ (9)
Accuracy Reward: Our training data mainly contains five kinds of tasks: Video Temporal Grounding (VTG), Multiple Choice (MC), MC&VTG (GLUE), ReOrdering (RO), and Open-Ended (OE), illustrated in Fig. 3. Except for the OE task, we use rule matching. And for the OE task, we use an open-source verification model (VM) [xVerify]. Given the question $Q$, the ground truth answer $G$, and the model’s generated answer $A_{\text{gen}}$ as input, we obtain the token probabilities $P_{Ic}$ and $P_{C}$ for the Incorrect and Correct labels, respectively. Thus, the accuracy reward $R_{acc}$ is following:
$R_{a}(y)=\begin{cases}IoU(G,y),&\text{if VTG}\\ Match(G,y),&\text{if MC }\\ \frac{P_{C}}{P_{Ic}+P_{C}},&\text{if OE}\\ Match(G,y)+IoU(G,y),&\text{else}\end{cases}$ (10)
where $Match$ is defined as choosing the same options or having the same sorting order.
Anti-hallucination Reward: In sec. 3, Factual-Aware Evaluator($E$) is trained for anti-hallucination reward. The reasoning model will output multiple visual evidences to describe video clips. Each evidence contains three pieces of information: $t_{start}$, $t_{end}$, and descdesc. The probability that $E$ judges whether the description is correct is:
$P_{Y/N}=E\left(clip(Video,t_{start},t_{end}),desc\right),$ (11)
where clip is defined as using ffmpeg to obtain video clips, and $Y$ means the description is correct. We also need to consider the quantity of evidence and ensure that different pieces of evidence do not describe the same content. For this purpose, the attenuation coefficient $w$ is defined as the product of the maximum IoU and ROUGE between the current evidence and other evidence:
$w_{i}=1-\max_{j\neq i}\left(\text{IoU}(e_{i},e_{j})\times\text{ROUGE}(e_{i},e_{j})\right).$ (12)
Thus, our anti-hallucination reward ($R_{h}$) is defined as a weighted sum of evidence scores, normalized by a non-linear function of $n$:
$R_{h}(y)=\frac{1}{max(0.6+0.8\times n,n)}\sum_{i=1}^{n}\frac{w_{i}P_{Y,i}}{P_{N,i}+P_{Y,i}}$ (13)
$R_{h}$ is only provided for the output where $R_{a}$ is greater than 0.5. Considering that the $R_{a}$ for the VTG and GLUE tasks is dense, $R_{h}$ may affect the IoU results.
The overall reward $R$ for a sample is computed as follows:
$R=R_{a}+0.5R_{f_{t}}+0.5R_{f_{e}}+0.2R_{h}$ (14)
Table 1: The results of Video Reasoning and General Benchmarks. For VCR-Bench, Video-$t$, and TVG-Bench, we re-tested the performance metrics of each model. The metric of TVG-Bench is R1@0.5. Considering the fairness, for Videorchat-R1.5, we limited the maximum number of frames to 64, the maximum number of pixels to $6000\times 28\times 28$, and used single inference mode. The other models were evaluated using the default settings in the paper. In the Training Data column, we compared the data scales in the cold start and RL phases, in the form of cold start + RL phase.
Model Training Data $\downarrow$ Reasoning General Temporal Avg
Video-Holmes MMVU(mc) VCR-Bench VSI-Bench Video-tt(mc) VideoMME TVG-Bench
Closed-Source Model
GPT-4o [openai2024gpt4ocard] - 42.0 67.4 46.9 34.0 46.6 71.9 - -
Gemini-2.0-Flash [comanici2025gemini] - 30.6 65.9 51.7 45.4 - - - -
3B Open-Source Model
Qwen2.5VL [baiQwen25VLTechnicalReport2025] - 17.0 53.0 24.8 28.6 36.9 52.1 14.3 32.4
Time-R1 [wang2025time] 0+2.5K 32.2 52.5 22.9 28.5 34.9 53.4 21.0 35.3
Video-RFT [wang2025videorft] 102+8K 34.3 52.6 25.8 31.9 35.2 54.7 10.6 34.7
Video-PLR (our) 38+6K 39.8 59.3 28.4 35.1 40.9 57.3 24.9 40.8
7B Open-Source Model
Qwen2.5VL [baiQwen25VLTechnicalReport2025] - 27.8 63.3 29.7 31.8 39.9 60.1 18.7 38.8
Time-R1 [wang2025time] 0+2.5K 37.1 64.2 29.6 28.4 39.0 60.9 29.6 41.3
Video-R1 [feng2025video] 165+8K 36.5 64.2 31.3 35.8 39.1 59.3 13.7 40.0
Video-RFT [wang2025videorft] 102+8K 39.8 68.5 30.6 36.8 40.3 59.8 15.4 41.6
Videochat-R1 [li2025videochat] 0+18K 33.0 63.8 30.0 31.3 39.1 60.8 30.1 41.2
Videochat-R1.5 [yan2025videochat] 0+80K 38.5 64.5 30.0 40.6 41.3 61.2 26.8 43.2
Video-PLR (our) 34+6K 42.6 64.7 32.4 38.3 42.7 61.8 31.7 44.9
Table 2: The results of hallucination benchmarks. is defined as rewriting questions as captions. ‘Diff’ refers to the difference in accuracy between the correct and incorrect categories. Compared with Qwen2.5VL, our FAE has better performance and achieves capabilities at the level of closed-source models.
Model VideoHallucer [wang2024videohallucer] HEAVEN [gao2025exploringhallucinationlargemultimodal] VidHalluc  [li2025vidhalluc]
[…]
GPT-4o [openai2024gpt4ocard] 74.6 0.9 - - 91.0
Gemini-1.5-Pro [geminiteam2024gemini15unlockingmultimodal] 62.9 41.3 - - 79.2
Qwen2.5VL [baiQwen25VLTechnicalReport2025] 67.8 22.4 63.7 28.9 89.2
FAE(ours) 75.4 (+7.5) 7.3 (-15.1) 72.9 (+9.2) 2.1 (-26.8) 90.3 (+1.1)
5 Experiment
5.1 Setups
Training Details For the Factual-Aware Evaluator, we use LoRA [hu2022lora] based on Qwen2.5VL 7B, training for one epoch with a learning rate of 1e-4, where $\lambda$ in formula 4.3 is 0.5, with a maximum of 16 frames and the total number of pixels limited to $1200\texttimes 28\texttimes 28$. For the reasoning model, training includes a cold-start SFT and RL using GRPO. The cold-start data incorporates 20K image CoT data from the Reason-RFT-CoT-Dataset [tan2025reason] and our VideoPLR-14K. We collected 6K questions for training, sourced from LLava-Video-178K [llava-video], Time-R1 [wang2025time], Cinepile [rawal2024cinepilelongvideoquestion], NextGQA [xiao2024can], and ANet [caba2015activitynet], as show in Fig. 3. Both SFT and RL, the learning rate is set to 2e-6, and train for one epoch. We set the number of generations to 8 and the batch size to 6, the $\beta$ eq. 4.3 is set to 0.01, and the maximum output length is set to 1600 tokens. The maximum number of frames and total pixels are set to $64$ and $2048\times 28\times 28$. Our training code is optimized based on ms-swift [zhao2024swiftascalablelightweightinfrastructure] to adapt to the anti-hallucination reward, which incurs significant computational latency. Thus, we implement GRPO training without additional time consumption.
Evaluation Details First, we evaluate the Factual-Aware Evaluator, including the VidHalluc [li2025vidhalluc], VideoHallucer [wang2024videohallucer], and HEAVEN [gao2025exploringhallucinationlargemultimodal]. Set the maximum number of frames to 16, the number of pixels to $2400\times 28\times 28$, and the temperature to 0. For video reasoning models, the evaluation benchmarks include three types: reasoning, general, and temporal. The reasoning benchmarks include: Video-Holmes [cheng2025videoholmesmllmthinklike], MMVU [zhao2025mmvu], VCR-benchmark [qi2025vcrbenchcomprehensiveevaluationframework],Video-tt [zhang2025towards], VSI-bench [yang2025thinking]. The general benchmark is VideoMME [fu2025video], and the temporal benchmark is TVG-Bench [wang2025time]. Set the maximum number of frames to 64, the number of pixels to $6000\times 28\times 28$, and the temperature to 0.
5.2 Hallucination Evaluation
The evaluation begins with the hallucination benchmark evaluation, as shown in Tab 2. We use VideoHallucer [wang2024videohallucer] for binary judgment, the open-source HEAVEN [gao2025exploringhallucinationlargemultimodal] data reformatted as captions, and the MCQ task from VidHalluc [li2025vidhalluc]. On VideoHallucer, FAE surpasses GPT-4o. It boosts accuracy over the base model by 7.5% and 9.2%. FAE improves judgment fairness, reducing bias towards ‘Y’ or ‘N’. It narrows the accuracy gap between correct and incorrect classifications by 15.1% and 26.8%. This reduction validates our data construction method and FAE’s reliability as an anti-hallucination reward, confirming the robustness and successful transfer of its learned abilities. Despite being trained exclusively on binary classification tasks, our model demonstrates strong generalization to the MCQ task. This indicates that the learned ability to judge hallucinations is robust and successfully transferred, validating the effectiveness of our training approach.
5.3 Main Results
We compared the models that were post-trained based on Qwen2.5VL. These baselines have different focuses in training. Among these models, Time-R1 and VideoChatR1 pay more attention to time perception, while VideoR1 and VideoRFT focus more on reasoning, and VideoChatR1.5 takes both into account. Our Video-PLR has the best performance and data efficiency at the same time.
SOTA Results: Our Video-PLR achieves a significant average improvement in both the 3B and 7B parameter scales, as shown in Tab. 1. In the comparisons with VideoHolmes, VCR-Bench, and Video-tt, our Video-PLR shows a stable improvement compared to VideoRFT and Video-R1 in terms of parameter scales of 3B and 7B. These evaluation benchmarks have more multi-hop problems, that is, the answers are distributed over multiple time periods and require joint reasoning. Our PLR paradigm has more advantages in handling such tasks. MMVU focuses more on scientific questions, and VideoRFT with 7B has the best performance, which may be because there is more domain knowledge in its training data. VideoChatR1.5 performs best in VSI-Bench. This benchmark requires a stronger spatial reasoning ability, and the training of VideoChatR1.5 includes a large number of image grounding and video tracking tasks, meeting this requirement. Surprisingly, our Video-PLR performs outstandingly in TVG-Bench. Time-R1 and VideoChat-R1 have trained a large number of VTG tasks, but do not have better performance than our model. This is related to their lack of a logic-rich cold start step and only optimizing the ability to output timestamps. The evaluation result of Video-PLR for the general benchmark, VideoMME, is also the best.
Data Efficient: While achieving the best performance, Video-PLR demonstrates excellent data efficiency. In the comparison at the 7B scale, our model only used a total of 34K cold starts and 6K for RL. The sampling for RL was 8 times in all cases, which can be approximately equivalent to the training consumption of 8 times the cold start. So, except for Time-R1, all the other models used far more training consumption than ours. Video-RFT and VideoChatR1 are approximately 2 times ours, Video-R1 is about 3 times ours, and VideoChat-R1.5 is close to 8 times ours. It should be noted that although we only used 0.5K of VTG training data, the effect has exceeded that of TimeR1 trained with 2.5K data in TVG-Bench.
5.4 Ablation Study
Table 3: Ablation study on the training data. It can be found that the combination of CoT and direct output can significantly enhance the hallucination evaluation ability.
Traing Set HAVEN  [li2025vidhalluc] VideoHallucer [wang2024videohallucer]
[…]
$\checkmark$ $\checkmark$ 72.9 2.1 75.3 7.3 54.2
Table 4: Ablate cold-start data and the hallucination reward design. We implement the same training configuration. Training with the anti-hallucination reward achieves the best performance.
Cold Start Rewards Video-Holmes MMVU VCR
w/o sft Only Acc 35.1 55.6 25.4
Video-R1(165K) Only Acc 37.8 58.6 27.2
Video-RFT(102K) Only Acc 37.6 54.4 27.6
+ Anti-H 37.0 58.2 28.5
our w/ PLR(34K) Only Acc 36.8 58.5 26.6
+ Anti-H w/o norm 36.9 56.0 26.2
+ Anti-H 39.8 59.3 28.4
Hallucination Training Data As shown in Tab. 3, the mixed training data helps the model achieve better performance in hallucination evaluation. Specifically, we provide the accuracy difference for the binary classification task. The addition of CoT data has significantly alleviated the classification bias. This proves the effectiveness of the CoT data design, forcing the model to answer based on the video content. Additionally, directly answering ‘Yes’ or ‘No’ is demonstrably more efficient than answering with CoT. The performance of the model trained with a mixture is also better than that trained only with CoT. The Gather column shows the comprehensive performance, which requires both answers to paired questions to be correct. Obviously, mixed training is superior to other solutions. Thus, Hybrid training yields more efficient responses and more accurate hallucination judgments.
Training Strategies As shown in Tab. 5, all the settings using cold start are better than those without cold start. Compared with Video-R1 and Video-RFT, the larger scale of the cold-start data brings better performance with only an accuracy reward. When adding the anti-hallucination reward (+Anti-H), our model achieved a more stable improvement compared to Video-RFT. This stems from the fact that the anti-hallucination at the segment level is more reliable, and the reliable perception stimulates the effectiveness of the PLR paradigm. Then we compare with the non-linear norm design of the anti-hallucination reward (eq. 13). The non-linear norm encourages the model to output more visual perception, because this design provides a lower reward for less visual evidence. When using the mean score of all evidence as the reward, the model tends to be conservative and only outputs reliable perceptual results, which creates a lack of evidence and leads to reduced performance. It should be noted that adding the anti-hallucination reward leads to a general improvement in the model, demonstrating the effectiveness of anti-hallucination reward design.
Supplementary materials provide more discussion, statistical data, training curves, and so on.
6 Conclusion
In this paper, we present Perception Loop Reasoning (PLR), a novel video reasoning paradigm that iteratively perceives and reasons about videos, ensuring sufficient visual evidence. To further ensure the reliability of visual perception, the Factual-Aware Evaluator participates in the reinforcement learning stage as an anti-hallucination reward, which encourages the model to provide sufficient and precise visual evidence. The hallucination preference data we carefully constructed enable the FAE to have reliable description-video consistency judgment, reaching the level of GPT-4o. Our Video-PLR achieves SOTA on both 3B and 7B parameter scales and is more data-efficient compared to previous methods. The PLR paradigm contributes a new implementation idea for video reasoning and proposes that perception is the basis of reasoning. We hope our contributions can promote the further development of video reasoning with an emphasis on the reliability of the reasoning process.
[…]
Supplementary Material
Overview
The supplementary material is organized as follows:
  • Sec. 7 discusses the impact of the anti-hallucination reward and outlines current limitations.
  • Sec. 8 presents additional experiments on reasoning honesty, training visualization, and inference efficiency analysis.
  • Sec. 9 details training hyperparameters and our parallel reward optimization strategy for efficiency.
  • Sec. 10 provides a comprehensive statistical analysis and bias quantification for the training datasets.
  • Sec. 111213 showcase qualitative data examples, detailed case studies, and data generation prompts.
7 Discussion
The anti-hallucination reward is one of the main contributions. In the main manuscript, the non-linear norm design and the anti-hallucination reward significantly affect the performance of the model, as illustrated in the ablation study. Sec. 8.2 further illustrates the trend changes in model outputs during training under different settings. The primary metrics observed are the changes in model output length and the quantity of evidence enumerated. The simplistic design of averaging anti-hallucination rewards has led to a decline in both output length and the number of perceptions, resulting in performance loss. This further supports that comprehensive perceptual evidence makes video reasoning more reliable and enhances performance. Also, the standard anti-hallucination encourages the model perception, resulting in an increase in both output length and the number of perceptions. Moreover, the crucial role of anti-hallucination reward is the Factual-Aware Evaluator (FAE). Although the hallucination benchmarks show the effectiveness of the FAE, the FAE is used with its probability. More detailed analysis and discussion are necessary to verify that FAE is superior to the baseline. Therefore, Sec. 8.5 presents the reward margins of different models when used as reward models, where a larger margin indicates a clearer reward signal.
Limitations. First, the PLR paradigm relies on the precise temporal grounding capabilities of the underlying VideoLLM. While some open-source models [chen2024internvl, chen2024expanding, zhu2025internvl3, zhang2024llavanext-video] currently face challenges in timestamp generation, we anticipate that rapid advancements in foundation models (e.g., Gemini series [geminiteam2024gemini15unlockingmultimodal, comanici2025gemini]) will alleviate this bottleneck. Second, due to computational constraints, we restricted our experiments to 3B and 7B parameter scales with efficient data sizes (VideoPLR-14K and AnetHallu-117K) and controlled resolutions, rather than training larger-scale (e.g., 72B) exhaustive models. Despite these constraints, Video-PLR achieves state-of-the-art performance and high data efficiency, confirming the validity of our paradigm.
8 More Reasoning Results
8.1 Honest Reasoning
Through an analysis of the training logs, we observed instances of unfaithful reasoning, a phenomenon also noted in recent literature [1]. Manifestations of such unfaithfulness include prematurely terminating the perception loop despite explicitly identifying a need for further evidence, as well as discrepancies between the reasoning chain’s conclusion and the final output answer. We posit that accurate perception is the prerequisite for sound logic; therefore, the primary objective of this work is to ensure the reliability of visual perception rather than the perfection of textual logic. Although optimizing textual logic falls outside the primary scope of this study, we have nonetheless included relevant experiments to address this aspect, as detailed in Tab. 5. Specifically, we employed Qwen2.5-32B-AWQ as a judge to evaluate the consistency between the generated reasoning and the final answer. The prompt instructs the model to assess solely whether the output is self-consistent, disregarding the presence of logical fallacies. The probability of the “Match” token is then utilized as the reward signal.
However, experimental results indicate that employing an off-the-shelf language model as a judge, even at the 32B parameter scale, fails to enhance overall performance. As shown in the last row of Tab. 5, the addition of the consistency reward (+ Honest) leads to a performance decline compared to the standard setting. While we observed a reduction in unfaithful behaviors during training, this did not translate into benchmark improvements. We hypothesize that this is due to the inherent preference bias of the non-fine-tuned judge model. The policy model likely overfitted to these specific preferences, resulting in rigid output patterns that hindered the reasoning flexibility required for complex video understanding tasks.
8.2 Training Visualization
Figure 5: This figure records the trend of changes in the number of tokens output by the model as training steps progress.
Figure 6: This figure records the trend of changes in the number of perceptions within the thinking by the model as training steps progress.
Table 5: Ablation Study Table from the main manuscript.
Cold Start Rewards Video-Holmes MMVU VCR
our w/ PLR(34K) Only Acc 36.8 58.5 26.6
+ Anti-H w/o norm 36.9 56.0 26.2
+ Anti-H 39.8 59.3 28.4
+ Anti-H +Honest 35.0 55.3 26.4
Thinking length and the number of perceptions are key indicators of model behavior. We find that the standard anti-hallucination reward encourages the model to generate more sufficient perceptions. Compared to the baseline without the anti-hallucination reward, the thinking length increases significantly, as shown in Fig. 5. This is attributed to the non-linear normalization strategy. Specifically, the reward degrades to a mean calculation when the perception count is at least three; however, for fewer perceptions, the reward is lower than the mean, acting as a penalty. Using the no-reward setting as a baseline, we observe that employing a simple mean reward increases thinking length but decreases the number of perceptions, as shown in Fig. 6. Tab. 5 indicates that the significant performance improvement is driven by the increased quantity of perceptions. The performance drop observed with longer outputs after removing the normalization further confirms that thinking length alone is not the key factor for performance gain.
Table 6: nextgqa
Model NextGQA
Acc mIoU
Qwen2.5-VL-7B 59.5 15.4
Videorchat-R1 70.6 32.4
Videorchat-R1-Thinking 69.2 36.1
Video-PLR (our) 77.9 31.1
8.3 NextGQA
Our RL training data includes 1k qa from NextGQA. Therefore, we evaluate our 7B Video-PLR and compare it with VideoChatR1, which is tuned with the full set of NextGQA. As shown in Tab. 6, VideoPLR surpasses all baselines in terms of accuracy. Specifically, it outperforms the Qwen2.5-VL-7B baseline by 18.4% and leads VideoChat-R1 by 7.3%. While VideoChat-R1-Thinking shows a slight advantage in mIoU, our model achieves the best overall answer accuracy with significantly less training data.
8.4 Reasoning Analysis
We evaluate the inference efficiency using VLLM on a single NVIDIA A40 GPU. As shown in Table 7, Video-PLR requires an average of 16.34s per query. While this latency is higher than the single-step baselines (e.g., 10.57s for Video-R1), it is a necessary trade-off for our dynamic and grounded reasoning, which is justified by the following observations.
First, regarding dynamic adaptability, shown in Fig. 7, Video-R1 exhibits a narrow distribution (100-500 tokens), suggesting a reliance on static, template-like responses. In contrast, Video-PLR shows a significantly broader dynamic range (300-1300 tokens). This indicates that our model adaptively adjusts its reasoning depth—allocating fewer tokens for simple queries and expanding its reasoning chain for complex tasks. Furthermore, regarding temporal awareness (Fig. 8), Video-PLR demonstrates a clear positive trend where both token count and evidence count increase with video length. This proves that the model is actively processing the temporal extent of the video rather than generating hallucinations of fixed length regardless of input complexity. Most importantly, as shown in Fig. 9, the increased computational cost translates directly into effective computation. There is a strong positive correlation between the generated token count and the number of valid evidence segments found. This confirms that the additional inference time is effectively utilized for the Perception Loop to gather necessary visual evidence, ensuring that the longer reasoning paths are grounded in facts.
Table 7: We report the average token length, token distribution range, and average inference latency per sample. All models are evaluated using VLLM on a single NVIDIA A40 GPU
Model Avg. Tokens Avg. Time (s)
Video-R1 298.13 10.57
Video-RFT 314.54 11.46
Video-PLR (Ours) 530.70 16.94
Figure 7: Compare the number of tokens output by Video R1 and VideoPLR in MMVU
Figure 8: Statistics of the number of tokens output by the model and the amount of evidence provided for videos of different lengths
Figure 9: The statistical analysis of the relationship between the number of tokens output by the model and the amount of evidence included shows a positive correlation.
8.5 Reward Margin Analysis
To validate the effectiveness of our FAE as a reward model, we evaluate its discriminative capability on the HAVEN dataset compared to baselines (SigLip and Qwen2.5-VL). As presented in Tab. 8, our model achieves a significantly higher AUC of 76.32%, surpassing Qwen2.5-VL and SigLip by 7.7% and 15.9%, respectively.
More importantly, we analyze the reward margin—the distinctiveness between correct and incorrect captions. As illustrated in the histograms (Fig. 12 vs. Fig. 11 and Fig. 10), our model exhibits a much sharper decision boundary compared to the baselines. The separation gap ($\Delta$) between the mean confidence for factual captions ($P_{yes}$) and hallucinations ($P_{no}$) reaches 29.62%, which is nearly $3\times$ that of SigLip (11.32%). This quantitative gap demonstrates that our FAE provides a more definitive and high-quality reward signal, effectively encouraging the policy to distinguish between hallucinations and facts during RL training, whereas baselines suffer from ambiguous boundaries that may confuse the policy.
Table 8: Quantitative comparison of reward distinctiveness on the HAVEN dataset. We report the AUC, the mean reward for factual ($R_{yes}$) and hallucinatory ($R_{no}$) captions, and the separation gap ($\Delta$). Our FAE provides the clearest decision boundary, offering a robust signal for RL training.
Model AUC Confidence Scores
Mean $R(\text{yes})$ $\uparrow$ Mean $R(\text{no})$ $\downarrow$ $\Delta$ (Gap) $\uparrow$
SigLip 60.46 40.96 29.64 11.32
Qwen2.5-VL 68.59 54.34 33.07 21.27
FAE (Ours) 76.32 65.44 35.82 29.62
Figure 10: Prediction margin distribution of SigLip.
Figure 11: Prediction margin distribution of Qwen2.5-VL.
Figure 12: Prediction margin distribution of our FAE.
9 Training Details
Figure 13: Comparison of the Standard MS-SWIFT GRPO Framework (left) and our Parallel Reward Optimization (right). In the standard framework, reward computation is a blocking operation. Our optimization decouples the reward computation into an asynchronous sub-thread. This allows the heavy inference of the Factual-Aware Evaluator to execute in parallel with the reference and policy model log-probability computations, effectively masking the external reward latency.
9.1 Training Efficiency Optimization
Efficiency Optimization for GRPO Training. Integrating the Factual-Aware Evaluator (FAE) as a reward model introduces significant computational overhead, as it requires additional inference passes for every generated sample. In the standard MS-SWIFT GRPO framework, the Compute Reward stage is sequential, meaning the training thread must idly wait for all reward models to return results before proceeding to calculating log probabilities (Get Refmodel Logps and Get Model Logps). This serial execution leads to a training bottleneck:
$T_{step}=T_{rollout}+T_{reward}+T_{logps}+T_{grad}$ (15)
To address this, we implemented a Parallel Reward Optimization strategy. As illustrated in the right panel of Fig. 13, we decouple the reward computation into an asynchronous sub-thread. Upon receiving the rollout data (Step_n Rollout), the main training thread immediately proceeds to the GPU-intensive log-probability calculations. Simultaneously, the sub-thread dispatches requests to the reward models. The main thread only performs an asynchronous wait (Async wait) after the log-probability calculations are complete. Since the time required for Get Refmodel Logps and Get Model Logps typically exceeds or matches the inference time of the reward models, the latency of the FAE is effectively masked. The optimized time consumption becomes:
$T_{step}=T_{rollout}+\max(T_{reward},T_{logps})+T_{grad}$ (16)
Deployment and Load Balancing. To ensure the reward computation throughput matches the training speed, we deployed the reward models on a separate dedicated inference cluster consisting of four NVIDIA RTX 4090 GPUs. Specifically:
  • The Verify Model (VM) for the Open-Ended task is deployed on 1 $\times$ RTX 4090.
  • The Factual-Aware Evaluator (FAE) is deployed on 3 $\times$ RTX 4090. We implemented a load-balancing mechanism that distributes the concurrent verification requests from the training batch across these three GPUs, ensuring optimal resource utilization and minimal latency.
This architecture allows us to utilize a complex, model-based anti-hallucination reward without increasing the overall wall-clock time of the training process.
9.2 Parameter Details
We present the parameter details of RL training in detail, as shown in Tab. 9 and Tab. 10.
Table 9: Training Setting of RL on 3B model
Parameter Value
DeepSpeed Zero3
Learning Rate $2\times 10^{-6}$
Batch Size 6
num_generations 6
temperature 1.0
top_p 0.95
top_k 50
$\beta$ 0.01
max_grad_norm 1
Base Model Qwen2.5VL 3B
Ref Model Qwen2.5VL 3B
Training Epochs 1
Max Frames 64
Max Pixels 1,505,280
Table 10: Training Setting of RL on 7B model
Parameter Value
DeepSpeed Zero3
Learning Rate $2\times 10^{-6}$
Batch Size 6
num_generations 8
temperature 1.0
top_p 0.95
top_k 50
$\beta$ 0.01
max_grad_norm 1
Base Model Qwen2.5VL 7B
Ref Model Qwen2.5VL 7B
Training Epochs 1
Max Frames 64
Max Pixels 1,505,280
10 Dataset Details
10.1 Dataset Information
In this section, we provide a detailed statistical analysis of the training data, covering both the AnetHallu-117K used for the anti-hallucination reward and the VideoPLR-14K used for the cold start stage. To ensure the Factual-Aware Evaluator (FAE) can robustly identify various types of errors without bias, we constructed the AnetHallu-117K dataset with a diverse distribution of hallucination categories. As shown in Fig. 14, the dataset covers five distinct hallucination types to ensure comprehensive evaluation capabilities. The distribution is relatively balanced, with Attribute Modification (21.8%), Detail Conflation (21.3%), Quantity Modification (21.0%), and Action Substitution (20.2%) each accounting for approximately 20% of the data. Additionally, Temporal Reordering accounts for 15.6%, focusing specifically on the chronological logic of video events. This diversity is crucial for preventing the model from overfitting to specific error patterns.
Beyond type diversity, we strictly controlled the label distribution to mitigate potential bias towards Yes or No answers. As illustrated in Fig. 15, the ratio of Positive (Yes) to Negative (No) samples is 46.6% to 53.4%, achieving a near-balanced state. To further enhance the model’s reasoning ability while maintaining inference efficiency, we employed a hybrid response format. As shown in Fig. 16, 34.1% of the data requires Chain-of-Thought (CoT) reasoning, while 65.9% consists of direct answers. We also addressed linguistic shortcuts by analyzing vocabulary distributions. The word clouds in Fig. 17 and Fig. 18 demonstrate a diverse and overlapping vocabulary across positive and negative captions. This confirms that our Text Bias Elimination pipeline effectively reduces linguistic cues that might otherwise lead to spurious correlations, ensuring the model judges based on visual content rather than word frequency.
For the cold start stage, we utilized the VideoPLR-14K dataset to initialize the Perception Loop Reasoning (PLR) paradigm. As shown in Fig. 19, this dataset is primarily constructed from two sources: NextQA (62.5%) and ActivityNet (37.5%). NextQA provides rich causal and temporal logic questions, which were re-annotated with timestamped reasoning chains to train the model’s sequential perception ability. Complementing this, ActivityNet serves as the source for the ReOrdering task, focusing on non-sequential event sorting and dense captioning to enhance the model’s global understanding of video structures. Finally, we analyzed the distribution of token lengths within the generated Chain-of-Thought (CoT) sequences. As presented in Fig. 15, the reasoning chains exhibit a substantial length, with a mean of 502.53 tokens and a median of 474.00 tokens. The distribution follows a natural pattern with a long tail extending beyond 1,200 tokens, indicating that the dataset encompasses a wide range of reasoning complexities. This length distribution confirms that our cold-start data provides rich and detailed reasoning contexts rather than superficial descriptions, thereby effectively initializing the model’s capability for deep perception loop reasoning.
Figure 14: Distribution of the five hallucination types in the AnetHallu-117K dataset.
Figure 15: The balance statistics of Yes and No answers in the binary judgment tasks.
Figure 16: The ratio of Direct Answer versus Chain-of-Thought (CoT) reasoning in the mixed training data.
Figure 17: Word cloud visualization of high-frequency concepts in the negative (hallucinated) captions.
Figure 18: Word cloud visualization of the vocabulary distribution in the positive (ground-truth) captions.
Figure 19: Source composition of the Cold-Start SFT data, consisting of NextQA and ActivityNet (ANet).
Figure 20: Distribution of token lengths in the generated reasoning chains (CoT).
10.2 Quantifying Textual Bias
To rigorously evaluate the effectiveness of our debiasing approach, we introduce the Vocabulary Polarity Score based on the Log-Odds Ratio with Dirichlet smoothing. For a given word $w$, its polarity score is defined as:
$\text{Score}(w)=\log\left(\frac{\hat{P}(w|\mathcal{D}_{pos})}{\hat{P}(w|\mathcal{D}_{neg})}\right)$ (17)
where $\hat{P}(w|\mathcal{D}_{pos})$ and $\hat{P}(w|\mathcal{D}_{neg})$ represent the smoothed probabilities of word $w$ appearing in the positive and negative caption sets, respectively. We utilize Dirichlet smoothing (adding a pseudo-count of 1) to prevent numerical instability caused by zero-frequency words.
A score significantly greater than 0 indicates a strong association with positive captions, while a score significantly less than 0 implies a negative bias. To quantify the overall bias of the dataset, we define the Mean Absolute Polarization (MAP):
$\text{MAP}=\frac{1}{|V|}\sum_{w\in V}|\text{Score}(w)|$ (18)
where $V$ denotes the vocabulary set. A lower MAP value indicates that the vocabulary distribution is more balanced, with fewer words serving as discriminative shortcuts for the model.
10.3 Analysis of Debiasing Effectiveness
Based on the proposed vocabulary polarity score, we design an iterative filtering mechanism to progressively neutralize the dataset. The detailed procedure is outlined in Algorithm 1.
We conducted a quantitative analysis to verify that our iterative filtering method reduces lexical cues without compromising semantic richness.
Distribution Shift. As visually presented in Figure 21, the original dataset (depicted in red) exhibits a dispersive distribution with heavy tails. This high variance suggests a significant presence of highly polarized words (extreme Log-Odds scores), which potentially serve as spurious correlations for the model. In contrast, the debiased dataset (green) demonstrates a distinct distributional shift towards a leptokurtic shape centered at zero. The contraction of the distribution tails confirms the successful removal of extreme lexical artifacts, resulting in a more neutral vocabulary profile.
Iterative Convergence. Furthermore, Figure 22 illustrates the dynamic optimization process of our method. The MAP score exhibits a monotonic downward trend over the course of 15 iterations. The steep decline in the initial phases (Iterations 0-6) corresponds to the elimination of the most obvious biased terms. As the process continues, the curve flattens, indicating that the algorithm progressively fine-tunes the dataset by addressing more subtle lexical imbalances. This convergence validates the stability of our penalty-based filtering strategy.
Figure 21: Comparison of Vocabulary Polarity Distributions. The plot contrasts the density of Log-Odds scores between the original dataset (Red) and the debiased dataset (Green).
Figure 22: Convergence of MAP Score. The line graph tracks the reduction in Mean Absolute Polarization across 15 filtering iterations.
11 Data Example
Figure 23: Visualization of the cold-start data in open-ended format.
Figure 24: Visualization of the cold-start data in multi-choice format.
Figure 25: Visualization of the hallucination preference data.
From Figure 23 to Figure 25, we present some cold-start data and hallucination preference data. The cold-start data is divided into two types: multi-choice and open-ended, as shown in Figure 23 and Figure 24, respectively.
12 Case Study
Figure 26: Case study on Video-Holmes. Video-PLR leverages its perception loop to capture dispersed visual cues (e.g., the social media usage and the nose touch). This mechanism enables the model to reason about the implicit causal relationships between events, successfully identifying the woman’s punitive intent where other models fail.
Figure 27: Case study on VCR-Bench. Although the video predominantly focuses on a single vehicle, a second car appears briefly (3s-10s). Compared to models that ignore these short segments, Video-PLR employs an iterative perception loop to conduct multi-pass fine-grained inspection, effectively mitigating critical information omission.
Here we present cases from Video-Holmes (Figure 26) and VCR-Bench (Figure 27). Video-Holmes is a bench that requires strong video inference capabilities. In this bench, Video-PLR demonstrates powerful inference capabilities, effectively utilizing its perception loop mechanism to perceive implicit event relationships in the environment. In contrast, VideoRFT and VideoChat-R1.5 remain at the level of understanding the scene itself, failing to capture the implicit relationships between events. In VCR-Bench, Video-PLR’s perceptual loop mechanism reduces the omission of key content and improves the model’s overall grasp of the video.
13 Prompts
In this section, we provide the detailed prompts used in our data construction pipeline. Fig. 28 to Fig. 31 illustrate the four-step pipeline for synthesizing the VideoPLR-14K cold-start data. Fig. 32 to Fig. 35 display the prompts for constructing the AnetHallu-117K dataset, including ground-truth generation, hallucination injection, and correction. Finally, Fig. 36 presents the inference system prompt for Video-PLR.
Figure 28: Step 1: Entity Name Extraction. The prompt extracts visual entities from the original QA pairs to standardize terminology for subsequent captioning.
Figure 29: Step 2: Frame Level Captioning. The prompt generates detailed descriptions for the current frame while maintaining object consistency with the previous frame.
Figure 30: Step 3: Initial CoT Generation. The prompt synthesizes frame-level descriptions into a structured reasoning chain with timestamps (Start, End, Description) and sufficiency assessment.
Figure 31: Step 4: CoT Logic Refinement. The prompt refines the initial chain of thought to ensure the logical flow is sound and explicitly excludes incorrect options based on visual evidence.
Figure 32: Detailed Factual Recaptioning. The prompt used to generate high-fidelity dense descriptions from video clips, serving as the ground truth ($C_{gt}$) for the hallucination dataset.
Figure 33: Negative Caption Generation. The prompt instructs the model to generate a specific type of hallucinated caption (e.g., Attribute Modification) based on the ground truth.
Figure 34: Positive Caption Generation (Correction). The prompt asks the model to correct a negative caption back into a factual description, ensuring paired sentence structures.
Figure 35: CoT Judgement Prompt. The prompt used for the Factual-Aware Evaluator (FAE) to determine whether a text description matches the video using a reasoning step.
Figure 36: Video-PLR Thinking Prompt. The inference prompt that instructs the model to use the <think> tag and perform the perception loop mechanism before answering.
Algorithm 1 Bias Eliminate Algorithm
1:
Input: $D_{in}$: Full dataset of (pos, neg) pairs
$N_{iter}$: Number of filter iterations
$P_{p},P_{n}$: Total filter percentage (pos, neg)
Output: $D_{out}$: The debiased dataset
[…]
4:for $i\leftarrow 1$ to $N_{iter}$ do
5:   $C_{p},C_{n}\leftarrow\textsc{GetAllCaptions}(D)$
6:   $M_{p},M_{n}\leftarrow\textsc{CreatePenaltyMaps}(C_{p},C_{n})$ $\triangleright$ Find bias
7:   $S_{p}\leftarrow\textsc{ScoreCaptions}(C_{p},M_{p})$ $\triangleright$ Score by bias
8:   $S_{n}\leftarrow\textsc{ScoreCaptions}(C_{n},M_{n})$
9:   $R_{p}\leftarrow\textsc{GetTopPercent}(S_{p},\delta_{p})$ $\triangleright$ Removal set
10:   $R_{n}\leftarrow\textsc{GetTopPercent}(S_{n},\delta_{n})$
11:   $D\leftarrow\textsc{FilterDataset}(D,R_{p}\cup R_{n})$ $\triangleright$ Filter
12:end for
[…]
15:return $D_{out}$

### v2

@@
 Extensive experiments show that our Video-PLR achieves the state-of-the-art in both 3B and 7B parameter scales and has the best data efficiency.
-Our code, models, and datasets are released on: https://github.com/BoweiPu/VideoPLR.
+Our code, datasets, and models are released on: https://github.com/BoweiPu/VideoPLR.
 Figure 1: Overview of VideoPLR.
@@
 1 Introduction
-Video Large Language Models (VideoLLMs) [llava-video, liLLaVAOneVisionEasyVisual2024, wangInternVideo2ScalingFoundation2024a] demonstrate remarkable potential in traditional video understanding.
+Video Large Language Models (VideoLLMs) [45, 21, 33] demonstrate remarkable potential in traditional video understanding.
 Building on this success, researchers are increasingly turning their attention to complex video tasks.
-Inspired by the Think-then-Answer paradigm [guo2025deepseek], many VideoLLMs [wang2025videorft, zhang2025tinyllava, feng2025video] tackle complex problems by generating intermediate reasoning steps.
+Inspired by the Think-then-Answer paradigm [17], many VideoLLMs [32, 43, 12] tackle complex problems by generating intermediate reasoning steps.
 This reasoning is based on video perception: the accurate description of visual information in the video.
@@
 Furthermore, anti-hallucination is the key to precise perception, which is the foundation of reasoning.
-However, previous models, including Open-source VideoLLMs [wang2024videohallucer, li2025vidhalluc] and vision-language models [zhai2023sigmoid], cannot provide reliable factual evaluation.
+However, previous models, including Open-source VideoLLMs [34, 22] and vision-language models [42], cannot provide reliable factual evaluation.
 To fill this gap, we introduce a Factual-Aware Evaluator (FAE), which is meticulously trained on our newly constructed AnetHallu-117K dataset.
-The ANetHallu-117K is constructed from the human-annotated ANet [caba2015activitynet] and contains 5 types of hallucination and hybrid CoT binary classification preference data, which can reduce discriminative bias.
+The ANetHallu-117K is constructed from the human-annotated ANet [10] and contains 5 types of hallucination and hybrid CoT binary classification preference data, which can reduce discriminative bias.
 FAE is comparable to GPT-4o in evaluating text-video consistency to serve as a reliable anti-hallucination reward for RL training.
@@
 Specifically, their constrained visual context simplifies grounding the descriptions, making anti-hallucination more readily identifiable.
-These two efforts jointly constitute our final model, Video-PLR via reinforcement learning [guo2025deepseek], achieving high data efficiency.
+These two efforts jointly constitute our final model, Video-PLR via reinforcement learning [17], achieving high data efficiency.
 For the RL stage, we utilize only 6K of mixed data, including five types of tasks.
@@
 Multimodal Chain-of-Thought (MCoT) simulates human reasoning by decomposing complex problems into intermediate steps, thereby enhancing transparency and performance.
-LLaVA-CoT [xu2024llava] is the first to implement a multimodal O1-like [openai2024o1] reasoning approach.
-Similarly, Video-of-Thought [feivideo] proposes a comprehensive five-stage framework.
-VideoEspresso [han2025videoespresso] and VideoCoT [wang2024videocot] achieve MCoT by invoking external tools as a multi-stage approach.
-Chain-of-Frames [ghazanfari2025chain] further integrates this capability into the model’s internal thinking process, enabling it to describe keyframe content.
+LLaVA-CoT [39] is the first to implement a multimodal O1-like [25] reasoning approach.
+Similarly, Video-of-Thought [11] proposes a comprehensive five-stage framework.
+VideoEspresso [18] and VideoCoT [35] achieve MCoT by invoking external tools as a multi-stage approach.
+Chain-of-Frames [16] further integrates this capability into the model’s internal thinking process, enabling it to describe keyframe content.
 However, single-frame descriptions are static, making it difficult to perform dynamic analysis across frames.
@@
 2.2 Reasoning VideoLLMs
-VideoR1 [feng2025video] first achieves R1-like [guo2025deepseek] video reasoning and utilizes GRPO [shao2024deepseekmath] for training.
-However, its cold-start data is sampled from Qwen2.5VL 72B [bai2025qwen25vltechnicalreport].
+VideoR1 [12] first achieves R1-like [17] video reasoning and utilizes GRPO [29] for training.
+However, its cold-start data is sampled from Qwen2.5VL 72B [1].
 This strategy overlooks the fact that even models with larger parameter counts lack video reasoning training.
-To enhance video perception, VideoR1 and DeepVideo-R1 [park2025deepvideo] enhance the model’s perception by modifying the standard GRPO.
-VideoRFT [wang2025videorft] advances this further by restructuring the CoT into a fixed pipeline: task, perception, reasoning, and conclusion.
-It utilizes SigLip [tschannen2025siglip] during the perception stage to mitigate hallucinations.
-TimeR1 [wang2025time] and VideoChatR1 [li2025videochat] place a greater emphasis on temporal perception and often lack consideration for general-purpose tasks.
-VideoChatR1.5 [yan2025videochat] achieves tool-use for video trimming based on temporal reasoning.
-ARC-HunyuanVideo [ge2025archunyuanvideo7bstructuredvideocomprehension] can generate dense video descriptions with timestamps, but lacks the ability to reason within these segments, and also does not address the hallucination problem.
+To enhance video perception, VideoR1 and DeepVideo-R1 [26] enhance the model’s perception by modifying the standard GRPO.
+VideoRFT [32] advances this further by restructuring the CoT into a fixed pipeline: task, perception, reasoning, and conclusion.
+It utilizes SigLip [31] during the perception stage to mitigate hallucinations.
+TimeR1 [36] and VideoChatR1 [23] place a greater emphasis on temporal perception and often lack consideration for general-purpose tasks.
+VideoChatR1.5 [40] achieves tool-use for video trimming based on temporal reasoning.
+ARC-HunyuanVideo [15] can generate dense video descriptions with timestamps, but lacks the ability to reason within these segments, and also does not address the hallucination problem.
 In contrast, our reasoning paradigm integrates temporal perception and execution, anti-hallucination training, and cyclical evidence gathering, representing a more advanced approach to reasoning training.
 2.3 Video Hallucination
-Longer reasoning chains reduce attention to visual input, thereby leading to hallucinations [liu2025more].
+Longer reasoning chains reduce attention to visual input, thereby leading to hallucinations [24].
 Consequently, hallucination control is crucial for video models.
-VideoRFT [wang2025videorft] utilizes SigLip [tschannen2025siglip].
+VideoRFT [32] utilizes SigLip [31].
 Nevertheless, SigLip is not optimally suited for video.
@@
 However, the current models cannot achieve this.
-VidHalluc [li2025vidhalluc] evaluates hallucinations from three key dimensions, while VideoHallucer [wang2024videohallucer] focuses on the bias of judging the correctness of descriptions.
+VidHalluc [22] evaluates hallucinations from three key dimensions, while VideoHallucer [34] focuses on the bias of judging the correctness of descriptions.
 To address this, we provide a controllable data generation pipeline for hallucination pairs and introduce a video hallucination judgement model.
@@
 Figure 4: Comparison of caption lengths in ANetHallu-117K
-To provide reliable consistency between textual descriptions and visual content, Factual-Aware Evaluator is trained on AnetHallu-117K, which is based on human-annotated ActivityNet (ANet) [caba2015activitynet].
+To provide reliable consistency between textual descriptions and visual content, Factual-Aware Evaluator is trained on AnetHallu-117K, which is based on human-annotated ActivityNet (ANet) [10].
 In this section, we describe in detail the caption construction pipeline and the preference training pipeline, as shown at the top of Fig.
@@
 This mixed data strategy ensures the model not only judges the correctness of the caption but also learns why, thereby achieving robust and efficient video factual evaluation.
-Odds Ratio Preference Optimization (ORPO) [hong2024orpo] is employed for training.
+Odds Ratio Preference Optimization (ORPO) [19] is employed for training.
 The total loss function $\mathcal{L}_{ORPO}$ for ORPO is defined as a weighted sum of the SFT loss and the Odds Ratio (OR) preference loss:
@@
 To accommodate the reasoning for these two sequences, we have covered two types of tasks.
-For the sequential part, we select NextQA [xiao2021next], which contains causal and temporal action reasoning.
-For the disordered part, we set event ordering based on ActivityNet [caba2015activitynet] named ReOrdering.
+For the sequential part, we select NextQA [37], which contains causal and temporal action reasoning.
+For the disordered part, we set event ordering based on ActivityNet [10] named ReOrdering.
 The challenge is the lack of timestamp and reasoning annotations within NextQA.
@@
 4.3 Training with GRPO and Reward Design
-The Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] process begins by sampling $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$ for a given query $q$ and evaluating their rewards $\{r_{i}\}_{i=1}^{K}$.
+The Group Relative Policy Optimization (GRPO) [29] process begins by sampling $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$ for a given query $q$ and evaluating their rewards $\{r_{i}\}_{i=1}^{K}$.
 The rewards are then normalized to compute the advantage $A_{i}$ for each response:
@@
 Except for the OE task, we use rule matching.
-And for the OE task, we use an open-source verification model (VM) [xVerify].
+And for the OE task, we use an open-source verification model (VM) [3].
 Given the question $Q$, the ground truth answer $G$, and the model’s generated answer $A_{\text{gen}}$ as input, we obtain the token probabilities $P_{Ic}$ and $P_{C}$ for the Incorrect and Correct labels, respectively.
@@
 Closed-Source Model
-GPT-4o [openai2024gpt4ocard] - 42.0 67.4 46.9 34.0 46.6 71.9 - -
-Gemini-2.0-Flash [comanici2025gemini] - 30.6 65.9 51.7 45.4 - - - -
+GPT-4o [9] - 42.0 67.4 46.9 34.0 46.6 71.9 - -
+Gemini-2.0-Flash [7] - 30.6 65.9 51.7 45.4 - - - -
 3B Open-Source Model
-Qwen2.5VL [baiQwen25VLTechnicalReport2025] - 17.0 53.0 24.8 28.6 36.9 52.1 14.3 32.4
-Time-R1 [wang2025time] 0+2.5K 32.2 52.5 22.9 28.5 34.9 53.4 21.0 35.3
-Video-RFT [wang2025videorft] 102+8K 34.3 52.6 25.8 31.9 35.2 54.7 10.6 34.7
+Qwen2.5VL [2] - 17.0 53.0 24.8 28.6 36.9 52.1 14.3 32.4
+Time-R1 [36] 0+2.5K 32.2 52.5 22.9 28.5 34.9 53.4 21.0 35.3
+Video-RFT [32] 102+8K 34.3 52.6 25.8 31.9 35.2 54.7 10.6 34.7
 Video-PLR (our) 38+6K 39.8 59.3 28.4 35.1 40.9 57.3 24.9 40.8
 7B Open-Source Model
-Qwen2.5VL [baiQwen25VLTechnicalReport2025] - 27.8 63.3 29.7 31.8 39.9 60.1 18.7 38.8
-Time-R1 [wang2025time] 0+2.5K 37.1 64.2 29.6 28.4 39.0 60.9 29.6 41.3
-Video-R1 [feng2025video] 165+8K 36.5 64.2 31.3 35.8 39.1 59.3 13.7 40.0
-Video-RFT [wang2025videorft] 102+8K 39.8 68.5 30.6 36.8 40.3 59.8 15.4 41.6
-Videochat-R1 [li2025videochat] 0+18K 33.0 63.8 30.0 31.3 39.1 60.8 30.1 41.2
-Videochat-R1.5 [yan2025videochat] 0+80K 38.5 64.5 30.0 40.6 41.3 61.2 26.8 43.2
+Qwen2.5VL [2] - 27.8 63.3 29.7 31.8 39.9 60.1 18.7 38.8
+Time-R1 [36] 0+2.5K 37.1 64.2 29.6 28.4 39.0 60.9 29.6 41.3
+Video-R1 [12] 165+8K 36.5 64.2 31.3 35.8 39.1 59.3 13.7 40.0
+Video-RFT [32] 102+8K 39.8 68.5 30.6 36.8 40.3 59.8 15.4 41.6
+Videochat-R1 [23] 0+18K 33.0 63.8 30.0 31.3 39.1 60.8 30.1 41.2
+Videochat-R1.5 [40] 0+80K 38.5 64.5 30.0 40.6 41.3 61.2 26.8 43.2
 Video-PLR (our) 34+6K 42.6 64.7 32.4 38.3 42.7 61.8 31.7 44.9
@@
 Compared with Qwen2.5VL, our FAE has better performance and achieves capabilities at the level of closed-source models.
-Model VideoHallucer [wang2024videohallucer] HEAVEN [gao2025exploringhallucinationlargemultimodal] VidHalluc  [li2025vidhalluc]
+Model VideoHallucer [34] HEAVEN [14] VidHalluc  [22]
 […]
-GPT-4o [openai2024gpt4ocard] 74.6 0.9 - - 91.0
-Gemini-1.5-Pro [geminiteam2024gemini15unlockingmultimodal] 62.9 41.3 - - 79.2
-Qwen2.5VL [baiQwen25VLTechnicalReport2025] 67.8 22.4 63.7 28.9 89.2
+GPT-4o [9] 74.6 0.9 - - 91.0
+Gemini-1.5-Pro [8] 62.9 41.3 - - 79.2
+Qwen2.5VL [2] 67.8 22.4 63.7 28.9 89.2
 FAE(ours) 75.4 (+7.5) 7.3 (-15.1) 72.9 (+9.2) 2.1 (-26.8) 90.3 (+1.1)
@@
 5.1 Setups
-Training Details For the Factual-Aware Evaluator, we use LoRA [hu2022lora] based on Qwen2.5VL 7B, training for one epoch with a learning rate of 1e-4, where $\lambda$ in formula 4.3 is 0.5, with a maximum of 16 frames and the total number of pixels limited to $1200\texttimes 28\texttimes 28$.
+Training Details For the Factual-Aware Evaluator, we use LoRA [20] based on Qwen2.5VL 7B, training for one epoch with a learning rate of 1e-4, where $\lambda$ in formula 4.3 is 0.5, with a maximum of 16 frames and the total number of pixels limited to $1200\texttimes 28\texttimes 28$.
 For the reasoning model, training includes a cold-start SFT and RL using GRPO.
-The cold-start data incorporates 20K image CoT data from the Reason-RFT-CoT-Dataset [tan2025reason] and our VideoPLR-14K.
-We collected 6K questions for training, sourced from LLava-Video-178K [llava-video], Time-R1 [wang2025time], Cinepile [rawal2024cinepilelongvideoquestion], NextGQA [xiao2024can], and ANet [caba2015activitynet], as show in Fig.
+The cold-start data incorporates 20K image CoT data from the Reason-RFT-CoT-Dataset [30] and our VideoPLR-14K.
+We collected 6K questions for training, sourced from LLava-Video-178K [45], Time-R1 [36], Cinepile [28], NextGQA [38], and ANet [10], as show in Fig.
 3.
@@
 The maximum number of frames and total pixels are set to $64$ and $2048\times 28\times 28$.
-Our training code is optimized based on ms-swift [zhao2024swiftascalablelightweightinfrastructure] to adapt to the anti-hallucination reward, which incurs significant computational latency.
+Our training code is optimized based on ms-swift [47] to adapt to the anti-hallucination reward, which incurs significant computational latency.
 Thus, we implement GRPO training without additional time consumption.
-Evaluation Details First, we evaluate the Factual-Aware Evaluator, including the VidHalluc [li2025vidhalluc], VideoHallucer [wang2024videohallucer], and HEAVEN [gao2025exploringhallucinationlargemultimodal].
+Evaluation Details First, we evaluate the Factual-Aware Evaluator, including the VidHalluc [22], VideoHallucer [34], and HEAVEN [14].
 Set the maximum number of frames to 16, the number of pixels to $2400\times 28\times 28$, and the temperature to 0.
 For video reasoning models, the evaluation benchmarks include three types: reasoning, general, and temporal.
-The reasoning benchmarks include: Video-Holmes [cheng2025videoholmesmllmthinklike], MMVU [zhao2025mmvu], VCR-benchmark [qi2025vcrbenchcomprehensiveevaluationframework],Video-tt [zhang2025towards], VSI-bench [yang2025thinking].
-The general benchmark is VideoMME [fu2025video], and the temporal benchmark is TVG-Bench [wang2025time].
+The reasoning benchmarks include: Video-Holmes [6], MMVU [48], VCR-benchmark [27],Video-tt [46], VSI-bench [41].
+The general benchmark is VideoMME [13], and the temporal benchmark is TVG-Bench [36].
 Set the maximum number of frames to 64, the number of pixels to $6000\times 28\times 28$, and the temperature to 0.
@@
 The evaluation begins with the hallucination benchmark evaluation, as shown in Tab 2.
-We use VideoHallucer [wang2024videohallucer] for binary judgment, the open-source HEAVEN [gao2025exploringhallucinationlargemultimodal] data reformatted as captions, and the MCQ task from VidHalluc [li2025vidhalluc].
+We use VideoHallucer [34] for binary judgment, the open-source HEAVEN [14] data reformatted as captions, and the MCQ task from VidHalluc [22].
 On VideoHallucer, FAE surpasses GPT-4o.
@@
 Table 3: Ablation study on the training data. It can be found that the combination of CoT and direct output can significantly enhance the hallucination evaluation ability.
-Traing Set HAVEN  [li2025vidhalluc] VideoHallucer [wang2024videohallucer]
+Traing Set HAVEN  [14] VideoHallucer [34]
 […]
@@
 We hope our contributions can promote the further development of video reasoning with an emphasis on the reliability of the reasoning process.
-[…]
 Supplementary Material
@@
 First, the PLR paradigm relies on the precise temporal grounding capabilities of the underlying VideoLLM.
-While some open-source models [chen2024internvl, chen2024expanding, zhu2025internvl3, zhang2024llavanext-video] currently face challenges in timestamp generation, we anticipate that rapid advancements in foundation models (e.g., Gemini series [geminiteam2024gemini15unlockingmultimodal, comanici2025gemini]) will alleviate this bottleneck.
+While some open-source models [5, 4, 49, 44] currently face challenges in timestamp generation, we anticipate that rapid advancements in foundation models (e.g., Gemini series [8, 7]) will alleviate this bottleneck.
 Second, due to computational constraints, we restricted our experiments to 3B and 7B parameter scales with efficient data sizes (VideoPLR-14K and AnetHallu-117K) and controlled resolutions, rather than training larger-scale (e.g., 72B) exhaustive models.